# multiMentalRoBERTa: A Fine-tuned Multiclass Classifier for Mental Health Disorder

### v1

@@ 6 Class (With Stress) 5 Class (Without Stress) 6 Class (With Stress) 5 Class (Without Stress)
TABLE VIII: Robustness Analysis of Multiclass Classifiers on eRisk 2025 Dataset [30]
@@
Since no benchmark dataset exists for evaluating robustness in our five-class and six-class multiclass settings, we conducted additional experiments by collapsing the problem into a binary classification of depression versus non-depression. While benchmark datasets such as eRisk exist for binary depression detection, they do not extend to multiclass scenarios. To fill this gap, we used the eRisk 2025 dataset (Task 2)[30] to test robustness, aggregating posts into binary labels. Because our models were originally fine-tuned for multiclass classification rather than binary tasks, performance was lower in this setting. Nonetheless, as shown in Table VIII, our multiMentalRoBERTa model consistently achieved stronger recall and F1 than fine-tuned MentalBERT, demonstrating superior robustness.
@@ VI Bias, Fairness & Safety
Mental health AI classification systems face technical and ethical challenges related to bias. As the American Psychological Association [31] noted, AI must be developed and deployed with attention to transparency, informed consent, and the mitigation of systematic biases that could worsen disparities in psychological care. Training data drawn from narrow platforms or self-reported labels may embed cultural and linguistic biases that reduce generalizability across diverse populations. To promote fairness, we collected data from multiple sources, including the Stress Annotated Dataset (SAD), Reddit corpora, and neutral forums (see Section III-A), ensuring broader representation and reducing overfitting to a single community. We also addressed ethical concerns (see section III-A) by applying de-identification and safeguards against re-identification, aligning the model with responsible mental health research practices.

### v2

@@ 6 Class (With Stress) 5 Class (Without Stress) 6 Class (With Stress) 5 Class (Without Stress)
TABLE VIII: Robustness Analysis of Multiclass Classifiers on eRisk 2025 Dataset [parapar2025erisk]
@@
Since no benchmark dataset exists for evaluating robustness in our five-class and six-class multiclass settings, we conducted additional experiments by collapsing the problem into a binary classification of depression versus non-depression. While benchmark datasets such as eRisk exist for binary depression detection, they do not extend to multiclass scenarios. To fill this gap, we used the eRisk 2025 dataset (Task 2)[parapar2025erisk] to test robustness, aggregating posts into binary labels. Because our models were originally fine-tuned for multiclass classification rather than binary tasks, performance was lower in this setting. Nonetheless, as shown in Table VIII, our multiMentalRoBERTa model consistently achieved stronger recall and F1 than fine-tuned MentalBERT, demonstrating superior robustness.
@@ VI Bias, Fairness & Safety
Mental health AI classification systems face technical and ethical challenges related to bias. As the American Psychological Association [American-Psychological-Association] noted, AI must be developed and deployed with attention to transparency, informed consent, and the mitigation of systematic biases that could worsen disparities in psychological care. Training data drawn from narrow platforms or self-reported labels may embed cultural and linguistic biases that reduce generalizability across diverse populations. To promote fairness, we collected data from multiple sources, including the Stress Annotated Dataset (SAD), Reddit corpora, and neutral forums (see Section III-A), ensuring broader representation and reducing overfitting to a single community. We also addressed ethical concerns (see section III-A) by applying de-identification and safeguards against re-identification, aligning the model with responsible mental health research practices.
@@ VII Conclusion
Acknowledgment
We gratefully acknowledge the support of the Northwestern Mutual Data Science Institute (NMDSI) for funding this project.